1) Solve the following ODE with IVP 2y + 6y - 8y = 0 y(0) = 4 y(0) = -1
Answer #1

Given differential equation ,  \small 2y''+6y'-8y=0

Let  y = e be a solution of the above differential equation .

Now ,  y = e

y = memp, y = memo

Substituting these values in the given differential equation we get ,

  \small 2m^{2}e^{mx}+6me^{mx}-8e^{mx}=0

→ (2m² + 6m - 8)eme = 0

\small \Rightarrow 2m^{2}+6m-8=0 , Since  \small e^{mx}\neq 0 .

\small \Rightarrow 2(m^{2}+3m-4)=0

\small \Rightarrow m^{2}+3m-4=0

\small \Rightarrow (m+4)(m-1)=0

\small \Rightarrow m=-4,1

So the solution of the given differential equation is ,

\small y(t)=c_{1}e^{-4t}+c_{2}e^{t} where \small c_{1},c_{2} are arbitrary constant and assuming  \small t as independent constant , if the independent variable is  \small x you can replace \small t by  \small x .

Now we will use initial conditions \small y(0)=4 and  \small y'(0)=-1 to find the value of the constants  \small c_{1},c_{2} .

\small y(0)=4

\small \Rightarrow c_{1}+c_{2}=4\small ......(i)

\small y'(0)=-1

\small \Rightarrow -4c_{1}+c_{2}=-1   \small ......(ii)

\small (i)-(ii) gives ,

\small 5c_{1}=5

\small \Rightarrow c_{1}=1

Substituting value of  \small c_{1} in  \small (i) we get ,

\small 1+c_{2}=4

\small \Rightarrow c_{2}=3

Substituting value of  \small c_{1},c_{2} the reqired solution of the given differential equation is ,

Answer :  \small y(t)=e^{-4t}+3e^{t} .

.

.

.

.

If you have doubt or need more clarification at any step please comment and don't forget to rate the answer . Your rating keeps us motivated .